OpenStudy (anonymous):

15x-10y=5 15y=-4+5x solve by substitution method with a solution set

OpenStudy (nottim):

ok, so rearrange one of the equations to solve for x or y

OpenStudy (anonymous):

y=1/15(5x-4)

OpenStudy (nottim):

now, sub that into the other equation (for y)

OpenStudy (anonymous):

15x-10(1/15(5x-4)=5 this is where iam having problems :(

OpenStudy (nottim):

what, coud you draw that out or use the equation editor?

OpenStudy (nottim):

i meant to say wait...no difference?

OpenStudy (anonymous):

\[15x-10(1/15(5x-4)=5\]

OpenStudy (nottim):

well, you know bedmas right? do it so in such steps

OpenStudy (anonymous):

 the thing i get confused when using fractions